# Latent Wavelet Diffusion: Enabling 4K Image Synthesis for Free

### v2

@@ 4.2 Quantitative Results
2K Results. Table 1 and the top block of Table 2 demonstrate consistent improvements from integrating LWD across multiple backbone models. On the HPD prompt dataset wu2023humanpreferencescorev2 , LWD reduces FID by up to 7% and LPIPS up to 6% while also achieving comparable MAN-IQA and QualiCLIP, indicating improved semantic alignment and perceptual quality. Moreover, on the Aesthetic dataset diffusion4k , these gains are observed across diverse architectures, reinforcing the generality and model-agnostic nature of our approach.
4K Results. On the Aesthetic-4K (the bottom block of Table 2) and HPD prompt dataset wu2023humanpreferencescorev2 (can be found in the supplementary materials), LWD outperforms baselines such as URAE and PixArt-Sigma, particularly in metrics like FID, CLIPScore and Aesthetics. The improvements are especially pronounced in regions with fine structural detail, such as hair, foliage, or architectural elements, highlighting LWD’s ability to scale effectively to ultra-high resolutions. These results suggest that frequency-aware supervision provides meaningful guidance even in challenging high-frequency regimes where baseline methods often struggle. LWD also achieves the highest GLCM score when paired with SD3-F16, and substantially improves Sana’s performance on both GLCM and compression ratio, indicating stronger fine-detail retention and texture fidelity, without compromising overall quality.
Table 2: Quantitative benchmarks of latent diffusion models on Aesthetic-Eval at $2048\times 2048$ and $4096\times 4096$. GLCM Score measures high-frequency texture richness using gray-level co-occurrence matrices, while Compression Ratio assesses visual complexity via JPEG file size heuristics, both introduced in Diffusion-4K (diffusion4k, ).
2K Model FID $\downarrow$ CLIPScore $\uparrow$ Aesthetics $\uparrow$ GLCM Score $\uparrow$ Compression Ratio $\downarrow$
SD3-F16 esser2024scaling 43.82 31.50 5.91 0.75 11.23
SD3-Diff4k-F16 diffusion4k 40.18 34.04 5.96 0.79 11.73
@@
PixArt-Sigma-XL pixartsigma 39.13 35.02 6.43 0.79 13.66
@@
4K SD3-F16 esser2024scaling - 33.12 5.97 0.73 11.97
SD3-Diff4k-F16 diffusion4k - 33.41 5.97 0.70 11.90
@@ 4.3 Qualitative Results
Zoomed-in comparisons highlight improved reconstruction of fine details (e.g., hair strands, eyelashes, small objects) that are often blurred or omitted by baselines. These results suggest that frequency-aware masking not only enhances local precision but does so in a context-sensitive manner, avoiding over-sharpening or artifacts. This indicates that LWD effectively balances fine detail refinement with structural coherence. More full-resolution results can be found in Appendix C and supplementary materials.
Figure 5: 4K images generated by LWD with different architectures.
@@ 4.4 Discussion
Limitations and Future Work. While LWD improves generation quality without architectural changes or inference overhead, it inherits certain limitations common to latent diffusion models. In particular, the compression imposed by the VAE can result in the loss of fine-grained semantic detail, potentially limiting performance in tasks that require precise spatial alignment or photorealistic accuracy.
@@ A.1 Discrete Wavelet Transform (DWT)
where $h_{\text{low}}$, $h_{\text{high}}$ are orthogonal wavelet filters (e.g., Haar or Daubechies), $\ast$ denotes convolution, and $\downarrow 2$ indicates downsampling by a factor of 2.
@@ A.2 Relevance Map Construction
This approach is theoretically supported by prior work in multiscale signal analysis, which demonstrates that the local maxima of wavelet detail coefficients correspond to structural singularities and perceptual boundaries [38]. While we do not explicitly track maxima across scales, aggregating the squared energies of the high-frequency subbands provides a lightweight approximation of local gradient magnitude, serving as an effective, interpretable proxy for visual saliency in latent space.
@@ Appendix C Additional Results for 4K
Upper caption: "Eiffel Tower was Made up of more than 2 million translucent straws to look like a cloud, with the bell tower at the top of the building, Michel installed huge foam-making machines in the forest to blow huge amounts of unpredictable wet clouds in the building’s classic architecture.". Lower caption: "Barbarian woman riding a red dragon, holding a broadsword, in gold armour."
To assess the efficacy of our proposed Latent Wavelet Diffusion (LWD), we conduct a detailed evaluation focusing on the challenging 4K resolution (4096$\times$4096).
@@ C.1 Quantitative Results
Table 3 presents the quantitative comparison on 4K image generation. We evaluate the generated images using several key metrics: MAN-IQA [67] and QualiCLIP [2], which assess perceptual image quality and alignment with textual prompts, respectively. Additionally, we compute the GLCM (Gray-Level Co-occurrence Matrix) score as a measure of texture complexity and detail richness in the generated high-resolution outputs. Finally, we report the Compression Ratio, indicating the compressibility of the generated images, which can be indicative of redundancy or lack of fine details.
@@ C.2 Qualitative Analysis
Table 3: Evaluation on ultra resolution (4096 $\times$ 4096) image generation task with [64] prompts.
Method MAN-IQA ($\uparrow$) QualiCLIP ($\uparrow$) GLCM Score $\uparrow$ Compression Ratio $\downarrow$ PixArt-Sigma-XL [9] 0.2935 0.2308 0.48 45.15 Sana-1.6B [65] 0.3288 0.4979 0.71 25.89 FLUX-1.dev [31] 0.3673 0.2564 - - URAE [69] 0.3850 0.3758 0.41 38.86 LWD + URAE 0.4011 0.3855 0.74 28.77
@@ Appendix D Frequency-Aware Evaluation
Upper caption: "An elderly man with a prominent, bushy beard and deep-set eyes wears a vibrant orange turban, his weathered face marked by lines of age and experience.". Lower caption: "A lone figure on a horse stands in a misty forest, gazing up at a tall, multi-tiered temple surrounded by towering trees and soft, diffused light. Steam rises from the rocks near a stream, creating an atmospheric scene of tranquility and mystery."
@@
Perceptual Metrics: For completeness, we also report FSIM [72] and MS-SSIM [62], which capture visual similarity and structural coherence, respectively. Both metrics are bounded between 0 and 1, with higher values indicating better perceptual quality.
@@
Upper caption: "A litter of golden retriever puppies playing in the snow. Their heads pop out of the snow, covered in.". Lower caption: "A curvy timber house near a sea, designed by Zaha Hadid, represent the image of a cold, modern architecture, at night, white lighting, highly detailed."
Table 4: Comparison of frequency-sensitive metrics across different methods on Aesthetic-4K [71] validation set.
@@
Upper caption: "A serene landscape features a winding river, flanked by trees with autumn foliage, leading to a rustic wooden cabin with a corrugated roof, set against a softly blurred background.". Lower caption: "A grand interior featuring intricate stained glass windows, an elaborate rose window, ornate frescoes depicting biblical scenes, and elegant chandeliers illuminating the richly decorated walls and arches."
@@ Appendix E Effect of Scale-Consistency Loss on Reconstruction Quality
Table 5: Quantitative reconstruction results of VAEs with and without the Scale-Consistency Loss 3.1 on Aesthetic-4K [71] validation set.

### v3

@@ 4.2 Quantitative Results
2K Results. Table 1 and the top block of Table 2 demonstrate consistent improvements from integrating LWD across multiple backbone models. On the HPD prompt dataset (Wu et al., 2023), LWD reduces FID by up to 7% and LPIPS up to 6% while also achieving comparable MAN-IQA and QualiCLIP, indicating improved semantic alignment and perceptual quality. Moreover, on the Aesthetic dataset (Zhang et al., 2025), these gains are observed across diverse architectures, reinforcing the generality and model-agnostic nature of our approach.
4K Results. On the Aesthetic-4K (the bottom block of Table 2) and HPD prompt dataset (Table 3), LWD outperforms baselines such as URAE and PixArt-Sigma, particularly in metrics like FID, CLIPScore and Aesthetics. The improvements are especially pronounced in regions with fine structural detail, such as hair, foliage, or architectural elements, highlighting LWD’s ability to scale effectively to ultra-high resolutions. These results suggest that frequency-aware supervision provides meaningful guidance even in challenging high-frequency regimes where baseline methods often struggle. LWD also achieves the highest GLCM score when paired with SD3-F16, and substantially improves Sana’s performance on both GLCM and compression ratio, indicating stronger fine-detail retention and texture fidelity, without compromising overall quality.
Figure 5: 4K images generated by LWD with different architectures.
Table 2: Quantitative benchmarks of latent diffusion models on Aesthetic-Eval at $2048\times 2048$ and $4096\times 4096$. GLCM Score measures high-frequency texture richness using gray-level co-occurrence matrices, while Compression Ratio assesses visual complexity via JPEG file size heuristics, both introduced in Diffusion-4K Zhang et al. (2025).
2K
Model FID $\downarrow$ CLIPScore $\uparrow$ Aesthetics $\uparrow$ GLCM Score $\uparrow$ Compression Ratio $\downarrow$
SD3-F16 Esser et al. (2024) 43.82 31.50 5.91 0.75 11.23
SD3-Diff4k-F16 Zhang et al. (2025) 40.18 34.04 5.96 0.79 11.73
@@
PixArt-Sigma-XL Chen et al. (2024) 39.13 35.02 6.43 0.79 13.66
@@
4K
SD3-F16 Esser et al. (2024) - 33.12 5.97 0.73 11.97
SD3-Diff4k-F16 Zhang et al. (2025) - 33.41 5.97 0.70 11.90
@@ 4.3 Qualitative Results
Zoomed-in comparisons highlight improved reconstruction of fine details (e.g., hair strands, eyelashes, small objects) that are often blurred or omitted by baselines. These results suggest that frequency-aware masking not only enhances local precision but does so in a context-sensitive manner, avoiding over-sharpening or artifacts. This indicates that LWD effectively balances fine detail refinement with structural coherence. More full-resolution results can be found in Appendix C.
@@ 4.4 Discussion
Limitations and Future Work. While LWD improves generation quality without architectural changes or inference overhead, it inherits limitations common to latent diffusion models. In particular, VAE compression can cause the loss of fine-grained semantic detail, potentially limiting performance in tasks requiring precise spatial alignment or photorealistic accuracy.
@@ 5 Conclusion
6 Reproducibility Statement
To ensure the reproducibility of our work, we have included the core Python script detailing our wavelet-based masking algorithm in the supplementary material. Further implementation details, including key hyperparameters and computational requirements, are also provided in the supplementary material. Following the review process, we will release our public GitHub repository, which will contain the complete implementation, training scripts, evaluation code, and the final pre-trained model checkpoints.
7 Ethical Statement
Our work is built upon publicly available datasets commonly used in the field of generative modeling. We acknowledge that these large-scale datasets may contain inherent societal biases, which our model could potentially learn and reproduce. We have used these datasets in accordance with their original licenses. Below, we discuss the potential societal impacts of our work.
7.1 Positive Impact
The primary goal of our research is to advance the state of high-resolution image generation, which has significant positive applications. These include empowering artists, designers, and content creators with more powerful creative tools; enhancing visual effects for entertainment and media; and potentially aiding in scientific visualization and data augmentation. By developing a method that improves quality with zero inference overhead, we aim to make high-fidelity generative AI more accessible and practical for a wider range of beneficial uses.
7.2 Potential Risks and Mitigation
We recognize that generative models can be misused for malicious purposes, such as creating misinformation (”deepfakes”), generating harmful or explicit content, and perpetuating societal biases. To mitigate these risks, we are committed to the following measures:
  1. 1.
    Responsible Release: We plan to release our code and models under a responsible AI license (e.g., a variant of the CreativeML Open RAIL-M license) that explicitly prohibits use for malicious, harmful, or unethical purposes.
  2. 2.
    Acknowledging Limitations: We are transparent about the limitations of our model and its potential to generate biased or factually incorrect content, as discussed in the main paper.
  3. 3.
    Encouraging Safe Deployment: We strongly encourage all downstream users of our models to implement their own safety filters, content moderation systems, and ethical guidelines before deploying any applications.
@@ A.1 Discrete Wavelet Transform (DWT)
where $_{\text{low}}, hh$$_{\text{high}}$ are orthogonal wavelet filters (e.g., Haar or Daubechies), $\ast$ denotes convolution, and $\downarrow 2$ indicates downsampling by a factor of 2.
@@ A.2 Relevance Map Construction
This approach is theoretically supported by prior work in multiscale signal analysis, which demonstrates that the local maxima of wavelet detail coefficients correspond to structural singularities and perceptual boundaries (Mallat & Hwang, 1992). While we do not explicitly track maxima across scales, aggregating the squared energies of the high-frequency subbands provides a lightweight approximation of local gradient magnitude, serving as an effective, interpretable proxy for visual saliency in latent space.
@@ Appendix C Additional Results for 4K
Upper caption: ”Eiffel Tower was Made up of more than 2 million translucent straws to look like a cloud, with the bell tower at the top of the building, Michel installed huge foam-making machines in the forest to blow huge amounts of unpredictable wet clouds in the building’s classic architecture.”. Lower caption: ”Barbarian woman riding a red dragon, holding a broadsword, in gold armour.”
To assess the efficacy of our proposed Latent Wavelet Diffusion (LWD), we conduct a detailed evaluation focusing on the challenging 4K resolution ($\times4096$).
@@ C.1 Quantitative Results
Table 3 presents the quantitative comparison on 4K image generation. We evaluate the generated images using several key metrics: MAN-IQA (Yang et al., 2022) and QualiCLIP (Agnolucci et al., 2024), which assess perceptual image quality and alignment with textual prompts, respectively. Additionally, we compute the GLCM (Gray-Level Co-occurrence Matrix) score as a measure of texture complexity and detail richness in the generated high-resolution outputs. Finally, we report the Compression Ratio, indicating the compressibility of the generated images, which can be indicative of redundancy or lack of fine details.
@@ C.2 Qualitative Analysis
Table 3: Evaluation on ultra resolution (4096 $\times$ 4096) image generation task with (Wu et al., 2023) prompts.
Method MAN-IQA ($\uparrow$) QualiCLIP ($\uparrow$) GLCM Score $\uparrow$ Compression Ratio $\downarrow$ PixArt-Sigma-XL (Chen et al., 2024) 0.2935 0.2308 0.48 45.15 Sana-1.6B (Xie et al., 2025) 0.3288 0.4979 0.71 25.89 FLUX-1.dev (Labs, 2024) 0.3673 0.2564 - - URAE (Yu et al., 2025) 0.3850 0.3758 0.41 38.86 LWD + URAE 0.4011 0.3855 0.74 28.77
Qualitative Results and Selection Protocol.
All qualitative examples shown in the paper were generated following a strict, reproducible protocol to prevent cherry-picking. Prompts were sourced directly from the original papers of the baseline models (e.g., URAE, Diffusion-4K) or the HPD benchmark dataset. For each prompt, the displayed image is the first output generated using a fixed random seed, applied identically to both the baseline and our LWD-enhanced model. We have included our code in the supplementary materials to ensure full transparency. While subjective preferences for certain images may vary, our method consistently improves objective indicators of texture and detail.
@@ Appendix D Frequency-Aware Evaluation
Upper caption: ”An elderly man with a prominent, bushy beard and deep-set eyes wears a vibrant orange turban, his weathered face marked by lines of age and experience.”. Lower caption: ”A lone figure on a horse stands in a misty forest, gazing up at a tall, multi-tiered temple surrounded by towering trees and soft, diffused light. Steam rises from the rocks near a stream, creating an atmospheric scene of tranquility and mystery.”
@@
Perceptual Metrics: For completeness, we also report FSIM (Zhang et al., 2011) and MS-SSIM (Wang et al., 2003), which capture visual similarity and structural coherence, respectively. Both metrics are bounded between 0 and 1, with higher values indicating better perceptual quality.
@@
Upper caption: ”A litter of golden retriever puppies playing in the snow. Their heads pop out of the snow, covered in.”. Lower caption: ”A curvy timber house near a sea, designed by Zaha Hadid, represent the image of a cold, modern architecture, at night, white lighting, highly detailed.”
Table 4: Comparison of frequency-sensitive metrics across different methods on Aesthetic-4K (Zhang et al., 2025) validation set.
@@
Upper caption: ”A serene landscape features a winding river, flanked by trees with autumn foliage, leading to a rustic wooden cabin with a corrugated roof, set against a softly blurred background.”. Lower caption: ”A grand interior featuring intricate stained glass windows, an elaborate rose window, ornate frescoes depicting biblical scenes, and elegant chandeliers illuminating the richly decorated walls and arches.”
@@ Appendix E Effect of Scale-Consistency Loss on Reconstruction Quality
Table 5: Quantitative reconstruction results of VAEs with and without the Scale-Consistency Loss 3.1 on Aesthetic-4K (Zhang et al., 2025) validation set.